OpenStudy (anonymous):

Satellites in orbits around Earth are: Outside of the reach of Earth gravity. Flying through the atmosphere similar to airplanes. Similar to high-speed projectiles. Always in circular orbits. @Abhisar

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (aayushi.somani):

option D.. :)

OpenStudy (anonymous):

wrong!

OpenStudy (aayushi.somani):

its right!

OpenStudy (anonymous):

nope its wrong :P

OpenStudy (aayushi.somani):

then what is the correct option? @Mashy

OpenStudy (anonymous):

i tagged u .. so that you can reason and answer it somy :P keplers first law rules out option D right? ..

OpenStudy (abhisar):

It will be C !

OpenStudy (abhisar):

The orbits can be elliptical too !

OpenStudy (aayushi.somani):

satellites in the orbit around earth cannot be compared to high speed projectile..!

OpenStudy (anonymous):

Abhisar is right :) why can't they? what is a projectile? @aayushi.somani

OpenStudy (aayushi.somani):

huh.. i might be wrong then.. i accept the apologies

OpenStudy (abhisar):

Projectiles which have the velocity = \(\sqrt{rg}\) will always stay in that orbit

OpenStudy (abhisar):

wr r is the distance of the orbit :)

OpenStudy (anonymous):

you accept apologies? xD.. lol u are hilarious!!!!! hahaha

OpenStudy (aayushi.somani):

i meant i apologise..! :P

OpenStudy (anonymous):

Don't apologies for being wrong :P.. being wrong is good.. it helps u learn.. we all just learnt something didn't we? ;-)

OpenStudy (abhisar):

Like i did just now !

OpenStudy (anonymous):

exacttlly @Abhisar ... ;-)

OpenStudy (aayushi.somani):

lol.. yeah.. thankq then @Mashy and @Abhisar

OpenStudy (abhisar):

;)

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)